OpenStudy (moonlight):

x^2 - 16/ x^2+4x+3 / x-4/x+1

OpenStudy (anonymous):

first flip the second one \[\frac{x^2-16}{x^2+4x+3}\times \frac{x+1}{x+4}\]

OpenStudy (anonymous):

ten factor and cancel \[\frac{(x-4)(x+4)(x+1)}{(x+3)(x+1)(x+4)}=\frac{x-4}{x+3}\]

OpenStudy (moonlight):

thanx

OpenStudy (anonymous):

yw